# ZeroComp: Zero-shot Object Compositing from Image Intrinsics via Diffusion

### v1

@@ Abstract
We present ZeroComp, an effective zero-shot 3D object compositing approach that does not require paired composite-scene images during training. Our method leverages ControlNet to condition from intrinsic images and combines it with a Stable Diffusion model to utilize its scene priors, together operating as an effective rendering engine. During training, ZeroComp uses intrinsic images based on geometry, albedo, and masked shading, all without the need for paired images of scenes with and without composite objects. Once trained, it seamlessly integrates virtual 3D objects into scenes, adjusting shading to create realistic composites. We developed a high-quality evaluation dataset and demonstrate that ZeroComp outperforms methods using explicit lighting estimations and generative techniques in quantitative and human perception benchmarks. Additionally, ZeroComp extends to real and outdoor image compositing, even when trained solely on synthetic indoor data, showcasing its effectiveness in image compositing.
@@ 2 Related work
Image compositing has a long history in vision and graphics. Of note, Reinhard et al. [55] studied color-based harmonization, later revisited in [52]. Issuing from the pioneering work by Burt and Adelson [9], seamless object blending has been studied in [49, 59, 67]. Lalonde et al. [41] inserted compatible objects from a large database into a target scene. Karsch et al.[33, 34] inserted synthetic 3D objects through the estimation of geometry and lighting annotations. Since then, image compositing techniques have approached the problem along different dimensions.
@@ 3.3 Preserving background fidelity
The ControlNet framework does not guarantee a perfect reconstruction of the background image. As demonstrated in previous studies, small details are susceptible to loss [93]. To mitigate this, we take inspiration from the differential compositing framework of Debevec [18] and generate a shadow opacity ratio from two predictions of ZeroComp. Denoting $f_{\theta}(\mathbf{i})$ as a full inference pass of the model on input intrinsic maps $\mathbf{i}=\{i_{d},i_{n},i_{a},i_{s}\}$ (containing depth, normals, albedo and shading resp., c.f. Sec. 3.1), we compute the shadow opacity ratio:
$R=\frac{f_{\theta}(\mathbf{i}_{\mathrm{comp}})}{f_{\theta}(\mathbf{i}_{\mathrm{bg}})}\,,$ (4)
where $\mathbf{i}_{\mathrm{comp}}$ and $\mathbf{i}_{\mathrm{bg}}$ are the intrinsic maps of the composite and background resp., see Eq. 2. Note that Eq. 4 is computed on grayscale and the result is clamped to $[0,1]$. We fix the diffusion with the same seed and use the same shading mask on the background $i_{s}$ to minimize discrepancies between both predictions. To further avoid unnecessary opacity unrelated to the object, we set the opacity to 1 if they are outside of the shading mask $m$ computed earlier (c.f. Sec. 3.2). A Gaussian blur with kernel $15\times 15$ and $\sigma=1.5$ is applied to $m$ to avoid blending artifacts. The final compositing equation is
@@ 6 Discussion
Acknowledgements
This research was supported by NSERC grants RGPIN 2020-04799 and ALLRP 586543-23, Mitacs and Depix. Computing resources were provided by the Digital Research Alliance of Canada. The authors thank Louis-Étienne Messier and Justine Giroux for their help as well as all members of the lab for discussions and proofreading help.

### v2

@@ Abstract
We present ZeroComp, an effective zero-shot 3D object compositing approach that does not require paired composite-scene images during training. Our method leverages ControlNet to condition from intrinsic images and combines it with a Stable Diffusion model to utilize its scene priors, together operating as an effective rendering engine. During training, ZeroComp uses intrinsic images based on geometry, albedo, and masked shading, all without the need for paired images of scenes with and without composite objects. Once trained, it seamlessly integrates virtual 3D objects into scenes, adjusting shading to create realistic composites. We develop a high-quality evaluation dataset and demonstrate that ZeroComp outperforms methods using explicit lighting estimations and generative techniques in quantitative and human perception benchmarks. Additionally, ZeroComp extends to real and outdoor image compositing, even when trained solely on synthetic indoor data, showcasing its effectiveness in image compositing.
@@ 2 Related work
Image compositing has a long history in vision and graphics. Of note, Reinhard et al. [55] study color-based harmonization, later revisited in [52]. Issuing from the pioneering work by Burt and Adelson [9], seamless object blending has been studied in [49, 59, 67]. Lalonde et al. [41] insert compatible objects from a large database into a target scene. Karsch et al.[33, 34] insert synthetic 3D objects through the estimation of geometry and lighting annotations. Since then, image compositing techniques have approached the problem along different dimensions.
@@ 3.3 Preserving background fidelity
The ControlNet framework does not guarantee a perfect reconstruction of the background image. As demonstrated in previous studies, small details are susceptible to loss [93]. To mitigate this, we take inspiration from the differential compositing framework of Debevec [18] and generate a shadow opacity ratio from two predictions of ZeroComp. Denoting $f_{\theta}(\mathbf{i})$ as a full inference pass of the model on input intrinsic maps $\mathbf{i}=\{i_{d},i_{n},i_{a},i_{s}\}$ (containing depth, normals, albedo and shading resp., c.f. Sec. 3.1), we compute the shadow opacity ratio $R=f_{\theta}(\mathbf{i}_{\mathrm{comp}})/f_{\theta}(\mathbf{i}_{\mathrm{bg}})$, where $\mathbf{i}_{\mathrm{comp}}$ and $\mathbf{i}_{\mathrm{bg}}$ are the intrinsic maps of the composite and background resp., see Eq. 2. Note that the ratio is computed on grayscale and the result is clamped to $[0,1]$. We fix the diffusion with the same seed and use the same shading mask on the background $i_{s}$ to minimize discrepancies between both predictions. To further avoid unnecessary opacity unrelated to the object, we set the opacity to 1 if they are outside of the shading mask $m$ computed earlier (c.f. Sec. 3.2). A Gaussian blur with kernel $15\times 15$ and $\sigma=1.5$ is applied to $m$ to avoid blending artifacts. The final compositing equation is
@@ 6 Discussion
Acknowledgments. This research was supported by NSERC grants RGPIN 2020-04799 and ALLRP 586543-23, Mitacs and Depix. Computing resources were provided by the Digital Research Alliance of Canada. The authors also thank L.-É. Messier, J. Giroux and members of the lab.